OpenStudy (anonymous):

Write an equation of a line in slope-intercept form with a slope of -4 and passing through the point (0, -3).

OpenStudy (texaschic101):

slope(m) = 4 (0,-3)...x = 0 and y = -3 y = mx + b -3 = 4(0) + b -3 = b equation is : y = 4x - 3